two trains each havin a speed of 80km/hr are headed towards each other on the same straight track ..a bird that can fly 60km/hr flies off the front of one train when they are 60 kms apart on reaching the other train heads for the first one again and so on ...what is the total distance travelled by the bird???

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

total time taken by train to meat = 60/(80+80)

                                                =3/8 hr

so in this tome interval distance moved by fly is = 60 * 3/8 =22.5 km

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin